Hi- Please help me with this word problem:
A rectangular field ha a perimeter of 118 feet. The width is 5 feet longer than half the length. Find the dimensions of the rectangle.
I have tried this so far:
P=118=2l+2w
w= 1/2+5
2l+2(1/2+5)=118
Tried it a few times, but not coming up with the correct answers. Please help. Thank you.

Answer by stanbon(75887)

A rectangular field ha a perimeter of 118 feet. The width is 5 feet longer than half the length. Find the dimensions of the rectangle.
I have tried this so far:
P=118=2l+2w
w= L/2+5
2L+2(L/2+5)=118
----
Divide thru by 2 to get:
L + (L/2) + 5 = 59
(3/2)L = 54
L = 36 (length)
---
solve for "W":
W = (L/2)+5
W = 23 (width)
==================
cheers,
Stan H.